### v1

# 5- and 6-membered rings: A natural orbital functional study

Abstract
The Global Natural Orbital Functional (GNOF) provides a straightforward approach to capture most electron correlation effects without needing perturbative corrections or limited active spaces selection. In this work, we evaluate both the original GNOF and its modified variant GNOFm on a set of twelve 5- and 6-membered molecular rings, systems characterized primarily by dynamic correlation. This reference set is vital as it comprises essential substructures of more complex molecules. We report complete-basis-set limit correlation energies for GNOF, GNOFm, and the benchmark CCSD(T) method. Across the Dunning basis sets, both functionals deliver a balanced and accurate description of the molecular set, with GNOFm showing small but systematic improvements while preserving the overall robustness of the original formulation. These results confirm the reliability of the GNOF family and its ability to capture dynamic correlation effects.
I Introduction
While the emergence of deep-learning and similar techniques has led to an improvement of parametrized methods, there is still room for ab-initio modern electronic structure methods. The latter are the unique alternative to practice discovery science, as recently shown by J. J. Eriksen et al.[1], who established the ground-state of Benzene by means of a blind challenge. However, emerging electronic structure methods require benchmarking, not only as a tool for comparison but a necessary test for validation. Benchmark studies provide a quantitative measure of the errors introduced by an approximation in computing different observables, which is essential for assessing the reliability of new approaches. Damour and co-workers [2] extended the aforementioned study of Benzene to a 12 molecular set compound by five- and six-membered rings. They investigated the performance and convergence properties of popular single-reference approaches, such as the Møller-Plesset perturbation series and the coupled-cluster (including iterative approximations) series, in comparison with full configuration interaction (FCI) correlation energy estimates. More importantly, the set included simple aromatic rings form the basis of more complex molecules of biological interest, so an accurate description is desired before going for larger and more complex systems. The motivation of the present study is to employ this molecular set to validate the performance of recent Natural Orbital Functional (NOF) approaches on molecules predominantly dynamic in correlation character.
NOF theory (NOFT),[3] as the one-particle reduced density matrix (1RDM) functional theory [4, 5, 6, 7, 8] in the natural orbital representation,[9, 10] along with other reduced density matrix methods,[11, 12] bridges the gap between DFT and wavefunction methods. Unlike the latter, which suffer from steep computational scaling, NOFT achieves a more efficient fifth-order scaling, reducible to fourth-order,[13] while accurately describing correlated electronic states. By utilizing the 1RDM and appropriately reconstructing the two-particle reduced density matrix (2RDM) from it, NOFT shows strong potential as a reliable alternative for multireference systems. Today, the complete active space self-consistent field (CASSCF)[14, 15] approach and its combination with second-order perturbation theory (CASPT2) [16, 17, 18, 19] remain the most reliable options. However, two major limitations significantly restrict the applicability of CASSCF and CASPT2: the need for active space selection and the high computational cost associated with a large number of strongly correlated orbitals. In contrast, NOF calculations correlate all electrons across all available orbitals within a given basis set, eliminating the complexities of active space selection. This makes NOFT particularly well-suited for problems such as bond-breaking and bond-forming reactions [20, 21], where a predefined active space may not be optimal. Additionally, the absence of user-defined input parameters removes arbitrariness and simplifies calculations, making NOFs more accessible to non-experts and appropriate to carry out studies without prior knowledge of the system, e.g. blind challenges.
Over the past two decades, NOFT has advanced significantly from both theoretical and computational perspectives. On the theoretical side, Piris and co-workers have developed a family of functionals known as PNOFs,[22, 23, 24, 25, 26] which continue to demonstrate their competitiveness with standard electronic structure methods. Their capabilities extend to various domains, including the description of excited states[27] and molecular dynamics,[28] as well as significant advancements in mitigating delocalization errors,[29] a persistent challenge in DFT. Additionally, PNOFs have contributed to understanding the ground-state spin state of iron(II) porphyrin,[30] a long-standing problem in electronic structure theory. More recently, NOFs have been employed for energy measurements on quantum computers, significantly improving efficiency within the variational quantum eigensolver (VQE) framework, giving rise to NOF-VQE.[31]
On the computational side, while NOFT calculations were initially constrained by high computational costs, recent advances have significantly improved their efficiency.[32, 33] A key development in this direction has been the incorporation of modern numerical techniques inspired by deep learning,[33] particularly momentum-based optimization methods such as the ADAM optimizer, which have accelerated the convergence of natural orbital calculations. These improvements have enabled NOFT to handle strongly correlated systems with up to 1000 electrons, the largest NOF calculations to date, making NOFT a viable tool for large-scale applications.
Despite these advances, NOFT remains underutilized, primarily due to two factors. First, NOF methods are not yet implemented in widely used electronic structure software packages. Although the open-source DoNOF program[34] for NOF calculations represents a significant step forward, broader integration is still needed. Second, accessible and systematic assessments of NOFs’ performance are scarce, making it difficult for researchers to gauge its reliability. In this vein, while the aforementioned GNOF approximation has been tested on strongly correlated models,[35, 36] its accuracy on systems dominated by dynamic correlation is undetermined yet, so a step forward in this direction is intended in the present work.
This article is organized as follows. The basics of NOFT are described in next section II, as well as the electron-pairing-based GNOF approximation and its modification GNOFm employed later on. In section III, the system set is introduced together with the methods that are used to compare with. Then, GNOF and GNOFm results are presented in section IV, together with reference CCSD(T) calculations. The article ends with a few remarks in section V.
II Electron-pairing-based NOFs
In this section, we outline the key concepts of NOFT to clarify its differences from commonly used approaches for studying strongly correlated systems. A more detailed description of NOFT and the approximations that define different NOFs can be found in Ref. [37]. Additionally, Ref. [38] presents a perspective on NOFT, discussing its fundamental concepts, strengths and weaknesses, current status, and potential future developments.
The energy of any NOF is typically expressed in terms of the set of NOs $\{\phi_{i}\}$ and their ONs $\{n_{i}\}$ as
$E[\mathrm{N},\{n_{i},\phi_{i}\}]=\sum_{i}n_{i}H_{ii}+\sum_{ijkl}D[n_{i},n_{j},n_{k},n_{l}]\langle ij|kl\rangle$ (1)
where the one- and two-electron integrals in the NO basis are given by
$H_{ii}=\int d{\bf r}\phi_{i}^{*}({\bf r})\left(-\frac{\nabla^{2}_{r}}{2}+v({\bf r})\right)\phi_{i}({\bf r})$ (2)
$\langle ij|kl\rangle=\int\int d{\bf r}_{1}d{\bf r}_{2}\frac{\phi^{*}_{i}({\bf r}_{1})\phi^{*}_{j}({\bf r}_{2})\phi_{k}({\bf r}_{1})\phi_{l}({\bf r}_{2})}{|{\bf r}_{2}-{\bf r}_{1}|}$ (3)
In Eq. (2), $v({\bf r})$ represents the nuclear potential determined by molecular geometry within the Born-Oppenheimer approximation, assuming no additional external fields. Unlike DFT, NOFT does not require a reconstruction for the one-electron part. However, the explicit form of the electron-electron interaction energy functional remains unknown, and different functional forms of $D[n_{i},n_{j},n_{k},n_{l}]$ lead to distinct NOFs.
The approximate functional (1) explicitly depends on the 2RDM,[39] requiring not only the N-representability of the 1RDM[40] but also that of the functional itself.[41] Specifically, the reconstructed $D[n_{i},n_{j},n_{k},n_{l}]$ must satisfy the same N-representability conditions as an unreconstructed 2RDM[42] to ensure the existence of a compatible N-electron system. Given their implicit dependence on the 2RDM, approximate functionals are best classified as NOFs rather than pure 1RDM functionals, as they are only defined in the NO representation.
In this article, we focus on electron-pairing-based functionals, which have proven particularly effective for describing strongly correlated systems and offer significant advantages from both theoretical and practical perspectives.[43] Accordingly, we consider $\mathrm{N_{I}}$ unpaired electrons that determine the system’s total spin $S$, while the remaining $\mathrm{N_{II}}=\mathrm{N-N_{I}}$ electrons form pairs with opposite spins, resulting in a net spin of zero for the $\mathrm{N_{II}}$ electrons.
We focus on the highest-multiplicity mixed state, where $2S+1=\mathrm{N_{I}}+1$ and the expectation value of $\hat{S}_{z}$ is zero. Consequently, the spin-restricted formalism can be applied, ensuring that all spatial orbitals $\{\varphi_{p}\}$ are doubly occupied within the ensemble and that $\alpha$ and $\beta$ spin particles have equal occupancies.[44]
Following the partitioning of electrons into $\mathrm{N_{I}}$ and $\mathrm{N_{II}}$, the orbital space $\Omega$ is divided into two subspaces: $\Omega=\Omega_{\mathrm{I}}\oplus\Omega_{\mathrm{II}}$. The subspace $\Omega_{\mathrm{II}}$ is composed of $\mathrm{N_{II}}/2$ mutually disjoint subspaces $\Omega{}_{g}$, each containing a reference orbital $\left|g\right\rangle$ for $g\leq\mathrm{N_{II}}/2$, along with $\mathrm{N}_{g}$ associated orbitals $\left|p\right\rangle$ for $p>\mathrm{N_{II}}/2$, formally expressed as
$\Omega_{g}=\left\{\left|g\right\rangle,\left|p_{1}\right\rangle,\left|p_{2}\right\rangle,...,\left|p_{\mathrm{N}_{g}}\right\rangle\right\}.$ (4)
Considering spin, the total occupancy of a given subspace $\Omega_{g}$ is 2, as expressed by the following pairing condition:
$\sum_{p\in\Omega_{g}}n_{p}=n_{g}+\sum_{i=1}^{\mathrm{N}_{g}}n_{p_{i}}=1,\quad g=1,2,...,\frac{\mathrm{N_{II}}}{2}.$ (5)
Similarly, $\Omega_{\mathrm{I}}$ consists of $\mathrm{N_{I}}$ mutually disjoint subspaces $\Omega_{g}$. Unlike $\Omega_{\mathrm{II}}$, each subspace $\Omega{}_{g}\in\Omega_{\mathrm{I}}$ contains only one orbital $g$ with an ON of $n_{g}=1/2$. Notably, each orbital holds a single electron, though its specific spin state, whether $\alpha$ or $\beta$, remains undetermined. From Eq. (5), it follows that the trace of the 1RDM equals the total number of electrons:
$2\sum_{p\in\Omega}n_{p}=2\sum_{p\in\Omega_{\mathrm{II}}}n_{p}+2\sum_{p\in\Omega_{\mathrm{I}}}n_{p}=\mathrm{N_{II}}+\mathrm{N_{I}}=\mathrm{\mathrm{N}}.$ (6)
The simplest electron-pair-based functional is PNOF5, which describes independent electron pairs,[45, 46] and its energy expression is given by:
$E\left[\mathrm{N},\left\{n_{p},\varphi_{p}\right\}\right]=E^{\text{intra}}+E_{\mathrm{HF}}^{\text{inter}}$ (7)
The intra-pair component is formed by summing the energies $E_{g}$ of electron pairs with opposite spins and the single-electron energies of unpaired electrons, specifically,
$E^{\mathrm{intra}}=\sum\limits_{g=1}^{\mathrm{N_{II}}/2}E_{g}+{\displaystyle\sum_{g=\mathrm{N_{II}}/2+1}^{\mathrm{N}_{\Omega}}}H_{gg}$ (8)
$E_{g}=2\sum\limits_{p\in\Omega_{g}}n_{p}H_{pp}+\sum\limits_{q,p\in\Omega_{g}}\Pi(n_{q},n_{p})L_{pq}$ (9)
where $L_{pq}=\left\langle pp|qq\right\rangle$ are the exchange-time-inversion integrals.[47] In Eq. (8), $\mathrm{\mathrm{N}_{\Omega}}=\mathrm{N_{II}}/2+\mathrm{N_{I}}$ denotes the total number of suspaces in $\Omega$. The matrix elements $\Pi(n_{q},n_{p})=c(n_{q})c(n_{p})$, where $c(n_{p})$ is defined by the square root of the ONs according to the following rule:
[…]
that is, the phase factor of $c(n_{p})$ is chosen to be +1+1 for the strongly occupied orbital of a given subspace $\Omega_{g}$, and 1-1 otherwise. The inter-subspace Hartree-Fock (HF) term is
$E_{\text{HF}}^{\text{inter}}=\sum\limits_{p,q}^{\mathrm{N}_{B}}\,{}^{\prime}\,n_{q}n_{p}\left(2J_{pq}-K_{pq}\right)$ (11)
where $J_{pq}=\left\langle pq|pq\right\rangle$ and $K_{pq}=\left\langle pq|qp\right\rangle$ are the Coulomb and exchange integrals, respectively. $\mathrm{N}_{B}$ denotes the number of basic functions considered. The prime in the summation indicates that only the inter-subspace terms are taken into account.
To enhance the inter-pair electron correlation, inter-subspace static and dynamic components must be added which lead to GNOF.[26] Its corresponding energy expression is given by:
$E\left[\mathrm{N},\left\{n_{p},\varphi_{p}\right\}\right]=E^{\mathrm{intra}}+E_{\mathrm{HF}}^{\mathrm{inter}}+E_{\mathrm{sta}}^{\mathrm{inter}}+E_{\mathrm{dyn}}^{\mathrm{inter}}\\$ (12)
where
$E_{\mathrm{sta}}^{\mathrm{inter}}=-\bigg(\sum_{p=1}^{N_{\Omega}}\sum_{q=N_{\Omega}+1}^{N_{\mathrm{B}}}+\sum_{p=N_{\Omega}+1}^{N_{\mathrm{B}}}\sum_{q=1}^{N_{\Omega}}+\sum_{p,q=N_{\Omega}+1}^{N_{\mathrm{B}}}\bigg)^{\prime}\Phi_{q}\Phi_{p}L_{pq}\\ -\frac{1}{2}\bigg(\sum_{p=1}^{N_{\text{II}}/2}\sum_{q=N_{\text{II}}/2+1}^{N_{\Omega}}+\sum_{p=N_{\text{II}}/2+1}^{N_{\Omega}}\sum_{q=1}^{N_{\text{II}}/2}\bigg)^{\prime}\Phi_{q}\Phi_{p}L_{pq}\\ -\sideset{}{{}^{\prime}}{\sum}_{p,q=N_{\text{II}}/2+1}^{N_{\Omega}}\Phi_{q}\Phi_{p}K_{pq}$ (13)
$E_{\text{dyn}}^{\text{inter}}=\sum\limits_{p,q=1}^{\mathrm{N}_{B}}{{}^{\prime\prime}\>}\left[\Pi(n_{q}^{d},n_{p}^{d})+n_{q}^{d}n_{p}^{d}\right]L_{pq}$ (14)
Here, $\Phi_{p}=\sqrt{n_{p}h_{p}}$ with $h_{p}=1-n_{p}$ being the hole. The second prime in Eq. (14) additionally excludes interactions between orbitals below the level $\mathrm{N_{II}}/2$. The dynamic contribution to the ON $n_{p}$ is defined as
$n_{p}^{d}=n_{p}\cdot e^{-\left(\dfrac{h_{g}}{h_{c}}\right)^{2}},\,\,p\in\Omega_{g}\,,\,\,g=1,2,...,\frac{\mathrm{N_{II}}}{2}.$ (15)
with $h_{c}=0.02\sqrt{2}$. The maximum value of $n_{p}^{d}$ is approximately 0.012, aligning with Pulay’s criterion, which states that an occupancy deviation of $\thickapprox 0.01$ from 1 or 0 is necessary for a NO to contribute to dynamic correlation.
Recently, a modified version of GNOF, denoted GNOFm, reintroduces the interactions between strongly occupied orbitals in the antiparallel spin blocks, as originally proposed in PNOF7.[24, 25] This refinement has shown improved accuracy for describing the singlet triplet energy gaps along the linear n-acene series.[33] Within this framework, the inter-subspace static component takes the following compact form:
$E_{\text{sta}}^{\text{inter}}=-\sum\limits_{p,q}^{\mathrm{N}_{B}}\,{}^{\prime}\Phi_{q}\Phi_{p}K_{pq}$ (16)
The solution is established by optimizing the energy with respect to the ONs and NOs, separately. Therefore, orbitals vary along the optimization process until the most favorable orbital interactions are found. All calculations have been carried out using the DoNOF code [34] and the recently implemented orbital optimization algorithm.[33]
III Motivation and Methodology
Comparisons between different NOFs are rare in the literature. Notable exceptions include studies on the behavior of various functionals, also beyond the electron-pairing approach, in the Hubbard Hamiltonian model [48, 49] and a rigorous assessment of 2RDM approximations that give rise to NOFs, evaluating their capacity to satisfy key properties of the exact functional.[50] Both comparative studies concluded that the functional N-representability is crucial for obtaining consistent results across different electronic correlation regimes. Consequently, we restrict our analysis to the electron-pairing-based NOFs presented in the previous section that enforce (2,2)-positivity conditions on the 2RDM. [42]
From a practical perspective, electron-pairing-based NOFs are particularly suited for describing strong correlation effects. In particular, the PNOF7 approximation was proven to be an efficient method for studying the Hubbard model and Hydrogen clusters described by a minimal basis set in one- and two-dimensions.[51, 52] Unfortunately, as recently shown by Lew-Yee and Piris,[33] PNOF7 could fail in molecular systems where dynamic correlation effects are non-negligible, and therefore the GNOF approximation is preferable for such systems. As briefly described in the previous section, GNOF aims to describe all electron correlation effects in a balanced manner, and numerous publications have demonstrated its ability to compete with standard electronic structure methods in different scenarios.[29, 30, 33, 26] Previous NOF approaches tried to retrieve dynamic correlation effects by terms of perturbation theory,[53, 24, 54] but including them into the functional itself gives access to correlated NOs and ONs. Nevertheless, while GNOF has been tested on model systems for strong correlation in one-, two- and three-dimensions,[35, 36] benchmarking its performance in systems dominated by dynamic electron correlation remains undone. In view of the results reported in Ref. [33], GNOF could be improved in complex correlation situations by a recent modification, so GNOFm is also included in the present work. This comparison, indeed, may help to clarify the delicate balance between dynamic and non-dynamic electron correlation terms in electron-pairing-based NOFs.
In Ref. [2], Damour et al. provided accurate FCI correlation energy estimates for twelve cases of five- and six-membered ring molecules, namely: Cyclopentadiene, Furan, Imidazole, Pyrrole, Thiophene, Benzene, Pyrazine, Pyridazine, Pyridine, Pyrimidine, s-Tetrazine, and s-Triazine. Hence, the set involves systems with atoms of the first to third lines of the periodic table. An schematic representation of the latter is shown in Fig. 1. In particular, Damour and co-workers reported optimized-orbital selected configuration interaction calculations for a correlation-consistent double-$\zeta$ Dunning basis set (cc-pVDZ),[55] as a reference for further studying the convergence of the Møller-Plesset perturbation theory series and the iterative approximate coupled-cluster series. Even in the context of the cc-pVDZ basis set, computing FCI result of these molecules is too computational demanding. Indeed, today carrying out coupled-cluster with singles, doubles, triples, and quadruples (CCSDTQ) calculations for molecules larger than benzene is prohibitively expensive or at least not practical.[1] This situation puts NOF approaches in an interesting position to run calculations employing larger basis sets from cc-pVDZ to cc-pV5Z. In the following, we use this molecular set to study GNOF and GNOFm correlation energies and their convergence with the size of the basis set. We provide complete-basis-set (CBS) estimates for these approximations, as well as for the ground-state gold standard coupled-cluster singles, doubles, and perturbative triples CCSD(T). The CCSD(T) calculations required a significant effort, especially in the largest cases at cc-pV5Z, which required around 1.4T of RAM, hence becoming infeasible for common computational configurations in contrast to NOF calculations. Following the work by Damour and co-workers, geometries of the molecular systems, obtained at the CC3/aug-cc-pVTZ level of theory, were extracted from Ref. [56].
Figure 1: 5- and 6-membered molecular rings studied along this work, as well as the corresponding numbering employed later on.
The DoNOF code [34] was employed for GNOF and GNOFm calculations, whereas CCSD(T) calculations were carried out with the PSI4 software package.[57] In contrast to Damour et al., no frozen core orbitals were considered in the present study. All electrons are correlated through all orbitals given in the basis set within the NOFT framework. The latter is, indeed, a strength of NOFs and their actual advantage with respect to typically used methods for multireference correlation, which require to define an active space where electrons are correlated. Finally, the resolution of identity approximation was used for integral evaluation in NOF calculations. [13] The latter was not employed in CCSD(T) calculations. However, as demonstrated[58] by DePrince III and Sherill, it would affect CCSD(T) energies, at most, in the order of a few $mE_{h}$, so in any case it alters neither the reported results nor the obtained conclusions.
IV Results
In this section, we analyze GNOF and GNOFm correlation energies for the aforementioned set of molecules, and compare them with CCSD(T) calculations. Note that correlation energies refer to the difference between energies given by a correlated method $E_{\mathrm{M}}$ and the Hartree-Fock energies $E_{\mathrm{HF}}$, i.e. $E_{\mathrm{corr}}=E_{\mathrm{M}}-E_{\mathrm{HF}}$.
Correlation energies for GNOF, GNOFm, and CCSD(T) are shown in Fig. 2 for increasing size correlation-consistent Dunning basis sets (cc-pVXZ, X=2-5). Here, molecules are ordered from smaller to larger correlation energies, according to the numbering presented in Fig. 1. Note that for thiophene (no. 3) with the cc-pV5Z basis, the reported values are not the raw results obtained directly from each method; the cc-pV5Z calculations show an unwarranted drop not observed for the other molecules or for thiophene with the remaining Dunning basis sets (cc-pVXZ, X=2-4). We attribute this behavior to a limitation in the sulfur cc-pV5Z basis set design. Consequently, we report an interpolated estimate from the X = 2, 3, and 4 Dunning sets. A detailed discussion of this issue is given in the second section of the Supplementary Material.
A first look to the plot reveals that NOF and CCSD(T) curves are roughly parallel, represented by dashed and solid lines, respectively, so the molecular description agrees for both methods independently of the size of the basis set, as well as of the different studied molecular rings. Probably the most noticeable disagreement is obtained for benzene (no. 6) when the GNOF/cc-pVTZ methodology is employed (dotted red curve). Correlation energy corresponding to the latter converges too rapidly when using GNOF in comparison with GNOFm and CCSD(T), thus the energy difference between cc-pVDZ and cc-pVTZ is slightly larger for GNOF than for the latter. Interestingly, this difference is removed when GNOFm is employed, which reveals a similar result to that obtained with CCSD(T). GNOFm does not provide a too low correlation energy for Benzene by using the cc-pVTZ basis set, in contrast with GNOF, so the behaviour obtained for this molecule follows the line of other systems in the set. Overall, GNOFm retrieves more correlation energy than its predecessor GNOF, and less than reference CCSD(T) calculations. Nevertheless, in the case of small basis sets there are a few exceptions where GNOFm and CCSD(T) compare very accurately or the former provides larger correlation energies than the latter, as it can be seen in Fig. 2.
Figure 2: Correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the selected set of molecules, obtained by using GNOF (dotted lines), GNOFm (dashed lines), and CCSD(T) (solid lines) with the $c$-pVXZ basis sets, $X=2,3,4,5$ being the cardinal number of the basis set. Molecules ordered according to the numbering given in Fig. 1.
Figure 3: Correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the selected set of molecules, obtained by using GNOF (dotted lines) and GNOFm (dashed lines) with the $c$-pVXZ basis sets ($X=2,3,4,5$), together with the resulting complete basis set (CBS) limit estimates. Details corresponding to the latter are given throughout the text, as well as in Fig. 4 and Table 1.
No. Systems GNOF GNOFm CCSD(T)
1 Cyclopentadiene -1083.4 -1114.4 -1156.5
2 Pyrrole -1103.5 -1153.9 -1197.9
3 Thiophene -1140.3 -1173.9 -1228.6
4 Furan -1150.0 -1182.9 -1225.9
5 Imidazole -1152.6 -1188.5 -1236.8
6 Benzene -1229.5 -1302.7 -1359.7
7 Pyridine -1304.5 -1342.5 -1397.8
8 Pyrazine -1333.0 -1378.4 -1439.1
9 Pyrimidine -1334.8 -1386.7 -1435.2
10 Pyridazine -1332.9 -1380.8 -1443.4
11 Triazine -1363.7 -1416.0 -1470.5
12 Tetrazine -1414.3 -1465.1 -1528.6
Table 1: Complete basis set (CBS) extrapolated correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the 12 molecular systems, computed using GNOF, GNOFm, and CCSD(T). An exponential extrapolation scheme, $E(X)=E_{\mathrm{CBS}}+a_{1}\cdot\exp(-a_{2}X)$, was employed with $X=2,3,4,5$ as the cardinal number of the basis set. For Thiophene, the extrapolation was performed using $X=2,3,4$.
As shown in Fig. 3, energy differences for the same molecule decrease as the size of the basis sets is augmented. This suggests that we are approaching the CBS limit for the reported correlation energies. Therefore, we computed the CBS limit correlation energies for the twelve molecules set using GNOF and GNOFm. An exponential function like extrapolation scheme was employed to obtain the CBS limit, $E(X)=E_{\mathrm{CBS}}+a_{1}\cdot\exp(-a_{2}X)$, $X=2,3,4,5$ being the cardinal number of the basis set. Interestingly, the form of the curve barely changes from cc-pVQZ to cc-pV5Z and from cc-pV5Z to the CBS limit, with the exception of the Benzene molecule in the case of GNOF calculations, which has been discussed before. Therefore, NOF calculations rapidly converge with the increasing size of Dunning correlation-consistent basis sets, so including more orbitals in the calculations just means lowering the total energy for GNOF and GNOFm calculations beyond cc-pVQZ. Previous studies [59] demonstrated similar results for different extrapolation schemes within the NOFT framework. Convergence of CCSD(T) correlation energies with the increasing size of the basis set can be seen in Fig. S1 from the Supplementary Material. Qualitatively, there are no significant differences with respect to the convergence obtained for GNOF and GNOFm.
The corresponding GNOF and GNOFm CBS estimated values are given in Table 1, which also presents extrapolations CCSD(T), performed using the same procedure. An inspection of CBS limit molecular correlation energies reveals an agreement within 100 mEh for GNOF in most cases, values that are even improved to around 50 mEh when GNOFm is utilized. The results shown in Table 1 are summarized in Fig. 4. GNOFm energies systematically get closer to CCSD(T) results when the static term between electron pairs is modified according to Eq. (16). In other words, Fig. 4 reveals that GNOFm CBS correlation energies reduce differences between GNOF and CCSD(T) to the half. Additionally, as it is shown in Figs. 2 and 3, the improvement is obtained for all basis sets studied.
Figure 4: Complete basis set (CBS) extrapolated correlation energies ($E-E_{\mathrm{HF}}$) in mEh for the 12 molecular systems, computed using GNOF, GNOFm, and CCSD(T). An exponential extrapolation scheme, $E(X)=E_{\mathrm{CBS}}+a_{1}\cdot\exp(-a_{2}X)$, was employed with $X=2,3,4,5$ as the cardinal number of the basis set. For Thiophene, the extrapolation was performed using $X=2,3,4$.
Finally, Fig. 4 presents a direct comparison of CBS-extrapolated correlation energies (in mEh) for the 12 rings obtained with GNOF, GNOFm, and CCSD(T). Both GNOF and its recent modification GNOFm provide very accurate descriptions of the five- and six-membered rings, as their traces lie close to and largely parallel with the CCSD(T) reference. Notably, Fig. 4 also shows a quantitative improvement of GNOFm over GNOF, consistent with Lew-Yee and Piris.[33] More importantly, the figure demonstrates that NOFs recover dynamic-correlation effects across the entire family of correlation-consistent Dunning basis sets and for all systems considered, proving the robustness of the Global NOF approach. Unlike earlier approximations, these functionals incorporate dynamic correlation within the energy expression itself and therefore do not require perturbative corrections.
V Closing remarks
This study assesses the performance of the most recent electron-pairing-based natural orbital functionals, GNOF and its modified variant GNOFm, on absolute correlation energies for five- and six-membered rings. This benchmark set, composed of simple aromatic rings of broad relevance, has previously been used to examine the performance and convergence properties of the Møller–Plesset series and coupled-cluster methods (including iterative approximations). Our results show that GNOFm attains quantitative agreement with the ground-state reference CCSD(T) across multiple sizes of the correlation-consistent Dunning basis sets. We also report complete-basis-set (CBS) extrapolated correlation energies for GNOF, GNOFm, and CCSD(T). A direct comparison between GNOFm and CCSD(T) indicates agreement to approximately 50 mEh, suggesting that the present data can serve as a useful benchmark for other quantum-chemistry approaches.
The error analysis and CBS extrapolations reported here for a representative set of five- and six-membered molecules clarify the capabilities and limitations of Global NOFs and help indicate when their application is most practical. While prior NOFT studies have often targeted systems with significant static correlation, the molecules investigated here are dominated by dynamic correlation. In this regime, GNOFm systematically improves upon GNOF in describing electron correlation, yielding a more balanced account of correlation effects.
The benchmarking dataset used here will be extended in a forthcoming work to guide refinements of the functional form, with particular emphasis on improving the treatment of electron correlation within the NOFT framework. Overall, the findings support the continued development of NOFs as a viable alternative to traditional density functionals and multireference wave-function methods. Their ability to capture both static and dynamic correlation without active-space selection makes them especially attractive for complex chemical systems, particularly when large molecules are involved. Future efforts should focus on further improving the balanced description of dynamic and non-dynamic correlations within NOFT to enhance accuracy across a broader range of chemical environments.
VI Acknowledgments
Financial support comes from the Eusko Jaurlaritza (Basque Government), Ref.: IT1584-22 and from the Grant No. PID 2021-126714NB-I00, funded by MCIN/AEI/10.13039/501100011033. J. F. H. Lew-Yee acknowledges the DIPC and the MCIN program “Severo Ochoa” under reference AEI/CEX2018-000867-S for post-doctoral funding (Ref.: 2023/74.) The authors acknowledge the technical and human support provided by both the DIPC Supercomputing Center and IZO-SGI (SGIker) of EHU and European funding (ERDF and ESF).